OpenStudy (anonymous):

simplify completely (x^2-100/x^2+2x-120)/(6 x+60/x-2)

OpenStudy (anonymous):

i'll give you a medal if you hurry up! :3

OpenStudy (ash2326):

We have \[ \frac{x^2-100}{x^2+2x-120}/\frac{6 x+60}{x-2}\] it can also be written as \[\frac{x^2-100}{x^2+2x-120}*\frac{x-2}{6 x+60}\] now a^2-b^2= (a+b)(a-b) so, (x^2-100)=(x+10)(x-10) let's find the factors of (x^2+2x-120)=(x+12)(x-10) so the express ion can be rewritten as \[\frac{(x-10)(x+10)}{(x+12)(x-10)}*\frac{x-2}{6 x+60}\] 6x+60=6(x+10) \[\frac{(x-10)(x+10)}{(x+12)(x-10)}*\frac{x-2}{6 (x+10)}\] let's cancel the common factors \[\frac{\cancel{(x-10)}\cancel{(x+10)}}{(x+12)\cancel{(x-10)}}*\frac{x-2}{6 \cancel{(x+10)}}\] so we get finally \[\frac{(x-2)}{6(x+12)}\]